need help with these 2 problems.Thanks for all your help.
write the equation of each root
1. 6, -2/3
2. -3/4, -8

Answer by ankor@dixie-net.com(22740)

need help with these 2 problems.Thanks for all your help.
write the equation of each root
:
1. 6, -2/3
Two factors
(x-6) = 0
and
(3x+2) = 0
FOIL
3x^2 - 18x + 2x - 12 = (3x^2 - 16x - 12) is the equation
:
2. -3/4, -8
Two factors:
(4x+3) = 0
and
(x+8) = 0
FOIL
4x^2 + 32x + 3x + 24 = (4x^2 + 35x +24) is the equation